What volume of hydrogen will be produced at stp by the reaction 55.5 grams of magnesium with excess water according to the following reaction

1
Expert's answer
2021-11-13T01:36:11-0500

Mg + 2H2"\rightarrow" Mg(OH)2 + H2

moles of Mg taken = (67.3/24) = 2.80 moles

Moles of H2 produce = 2.80moles

volume of H2 at STP = 2.80× 22.4 L = 62.8 L